# Organizing Unstructured Image Collections using Natural Language

### v4

@@ Abstract
Organizing unstructured image collections into semantic clusters is a long-standing challenge. Traditional deep clustering techniques address this by producing a single data partition, whereas multiple clustering methods uncover diverse alternative partitions, but only when users predefine the clustering criteria. Yet expecting users to specify such criteria a priori for large, unfamiliar datasets is unrealistic. In this work, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), which aims to automatically discover clustering criteria from large, unstructured image collections, revealing interpretable substructures without human input. Our framework, $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, treats text as a reasoning proxy: it concurrently scans the entire image collection, proposes candidate criteria in natural language, and groups images into meaningful clusters per criterion. To evaluate progress, we release COCO-4c and Food-4c benchmarks, each annotated with four grouping criteria. Experiments show that $\mathcal{X}$-Cluster effectively reveals meaningful partitions and enables downstream applications such as bias discovery and social media image popularity analysis. We will open-source code and data to encourage reproducibility and further research.
Organizing Unstructured Image Collections using Natural Language
Mingxuan Liu1    Zhun Zhong4    Jun Li6    Gianni Franchi3    Subhankar Roy5    Elisa Ricci1,2 1 University of Trento   2 Fondazione Bruno Kessler  3 ENSTA Paris, Institut Polytechnique de Paris 4 Hefei University of Technology   5 University of Bergamo   6 Technical University of Munich
Figure 1: Top: Open-ended Semantic Multiple Clustering (OpenSMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable clusters, under multiple shared themes or criteria, without requiring any prior knowledge. Bottom: Our system enables various applications like discovering novel biases in text-to-image (T2I) generative models. See our project webpage for code and data: here.
@@ 1 Introduction
When grouping a large, unstructured collection of unlabeled images, several questions arise: i) Which criterion should guide the grouping? ii) Do multiple, equally valid criteria exist? iii) Given a criterion, how can we uncover the corresponding semantic clusters? iv) What should the clusters’ granularity be? Because the answers depend on both the collection’s visual content and the objectives of a specific application, each question is inherently open-ended. Having the ability to automatically resolve them would unlock hidden insights in the data and enable valuable applications, such as social media recommendation systems (Cheng et al., 2024) and dataset auditing (Bianchi et al., 2023). footnotetext: † Corresponding author.
Previous studies have attempted to answer these questions. Deep Clustering (DC) approaches (Caron et al., 2018; Van Gansbeke et al., 2020) iteratively partition an unlabeled dataset but produce only a single partition. Multiple Clustering (MC) methods (Yu et al., 2024), instead, uncover several alternative clustering solutions, offering complementary perspectives. Yet both DC and MC ultimately reflect the inductive biases of the underlying networks, which can deviate from the partitions a user actually needs. Worse, the discovered criteria and clusters, denoted merely by index labels, lack human-interpretable semantics.
To give users explicit control over clustering, Text Criterion conditioned Multiple Clustering (TCMC) (Yao et al., 2024, 2025) approaches let the users specify the desired criterion in natural language e.g., “group playing cards by Suit.” TCMC methods (Kwon et al., 2024; Luo et al., 2024b; Yao et al., 2024, 2025) first employ large language models (LLMs) to generate candidate cluster semantics conditioned on the user-specified criterion, then use vision–language models to assign images to semantically related clusters. Although TCMC methods yield user-guided and interpretable partitions, they presume the user already knows valid grouping criteria for the entire collection. As datasets grow and scenes become more complex (e.g., millions of social-media images (Hu et al., 2014)) knowing such prior knowledge becomes impractical. This leads to the question: Can clustering criteria and semantic substructures (groups) be discovered directly from visual data, without any human input?
We formalize this challenge as the task of Open-ended Semantic Multiple Clustering (OpenSMC). As depicted in Figure 1(top), the objective in OpenSMC is to discover, in natural language, both the clustering criteria and their associated semantic clusters without relying on any human priors. The task is challenging because i) it requires concurrent reasoning over the entire image collection to surface valid criteria, posing a challenge to current vision systems, and ii) it is open-ended as neither user-defined knowledge about the clustering criteria nor clustering granularity is available. Tab. 1 contrasts OpenSMC with existing clustering paradigms. In addition, due to the lack of practical benchmarks with diverse criteria for evaluating the OpenSMC, we introduce two realistic and larger benchmarks–COCO-4c and Food-4c–with annotations for up to four criteria.
@@
To tackle OpenSMC, we leverage text as a compact proxy for reasoning because current vision models cannot yet encode and reliably reason over large image collections. Consequently, we propose $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a two-stage framework that combines multimodal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2025, 2023a) with text-only LLMs (Meta, 2024a). $\mathcal{X}$-Cluster comprises two modules: a Criteria Proposer, which jointly reasons over all images via their textual representation (embeddings, tags, or captions) to surface valid grouping criteria, and a Semantic Grouper, which uncovers semantic clusters and assigns images to them under each discovered criterion. As shown in Figure 1(top), $\mathcal{X}$-Cluster can help users explore unfamiliar image collections at scale, by automatically grouping the images into semantic clusters (e.g., “Surfing” and “Skateboarding”) under diverse, human-interpretable criteria (e.g., Activity). In this work, we investigate three designs for each module and demonstrate the flexibility and effectiveness of our framework.
Finally, we showcase both the practicality and versatility of $\mathcal{X}$-Cluster across several applications, including uncovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images. For instance, as shown in Figure 1(bottom), the method uncovers well-known biases (e.g., Gender) alongside less explored ones (e.g., Hair color) in DALL·E3 -generated images (Betker et al., 2023). These results confirm that $\mathcal{X}$-Cluster is an automated, practical tool for extracting interpretable insights from large-scale, unstructured visual data, opening new avenues for research and industry.
@@ 2 Related Work
Image clustering. Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self-supervision (Caron et al., 2018, 2020; Ren et al., 2024). Multiple clustering extends this idea, seeking multiple non-redundant partitions with data augmentations, diversity losses, or subspace methods (Qi and Davidson, 2009; Hu and Pei, 2018; Yao et al., 2023b; Metaxas et al., 2023; Yu et al., 2024). Despite steady progress, both paradigms share key limitations: i) their results reflect inductive biases of the network and training algorithm, hampering generalization beyond object-centric data and often misaligning with user intent; and ii) clusters are returned as numeric indices rather than human-readable names. In contrast, $\mathcal{X}$-Cluster derives both meaningful criteria and cluster names directly from unlabeled data.
@@
While $\mathcal{X}$-Cluster also uses text as its reasoning medium, it departs from TCMC systems in two key aspects: i) it automatically discovers the grouping criteria rather than relying on a user-supplied one; ii) it infers both the number and the names of clusters, without any user input. Together, these features make $\mathcal{X}$-Cluster fully automatic and scalable, and well suited to uncover novel, often overlooked insights in large image collections.
Topic Discovery. OpenSMC is also related to Topic Discovery (Blei et al., 2003) in NLP, which uncovers latent themes in large text corpora. Our work likewise seeks to reveal common thematic concepts, but from visual collections, and does so by using LLMs as the reasoning core.
3 OpenSMC
Task Definition: Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering (OpenSMC) is to build a system $\mathcal{H}$ that automatically i) discovers a set of $L$ grouping criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them. Formally, we define an OpenSMC system:
@@ 3 OpenSMC
An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language. By contrast, in the TCMC setting humans predefine the criterion $\mathcal{R}$ and the desired cluster count $K$.
Refer to caption(Prompts in App. E.1)(Prompts in App. E.2)
Figure 3: $\bm{\mathcal{X}}$-Cluster overview. (Left) The Criteria Proposer reason over the entire image set via their textual proxies and deposits diverse, natural language grouping criteria in a Criteria Pool. (Middle) The Semantic Grouper draws each criterion from the pool, discovers semantic clusters at three levels of granularity, and assigns images to their proper clusters. (Right) Aggregating these assignments reveals the collection’s semantic substructure across multiple criteria and granularities. We evaluate three design variants for each module and adopt the best-performing configuration (marked Refer to caption) as the default $\mathcal{X}$-Cluster. See App. E for implementation details.
Benchmark: Only a few benchmarks exist that are suitable for evaluating OpenSMC. As shown in Figure 2, current benchmarks: Fruit-2c (Muresan and Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), Clevr-4c (Vaze et al., 2024) are limited because they are object-centric (Fruit-2c), lacking photorealism (Clevr-4c), and too few criteria (e.g., at most three in Action-3c).
To alleviate this issue, we introduce two new benchmarks: Food-4c and COCO-4c. Examples of images are presented in Figure 2. We extended Food-101 dataset (Bossard et al., 2014) to create Food-4c which in total has four grouping criteria. Similarly, we extend COCO-val (Lin et al., 2014) to create COCO-4c that has four grouping criteria. We believe that the newly introduced more realistic benchmarks will play a critical role in evaluating OpenSMC systems. Further benchmark statistics and annotation details are provided in App. C.
4 Method
To recap, the objective of an OpenSMC system is twofold: i) discover meaningful grouping criteria, or themes, in an unstructured image collection, and ii) group the images into semantic clusters for each discovered criterion. This task is particularly challenging because it requires simultaneous reasoning over a large number of images, which the current vision systems are not equipped to handle.
System overview. We propose $\bm{\mathcal{X}}$-Cluster, a two-stage framework that is designed to perform joint reasoning over an entire image collection, as required by the OpenSMC task. At a high level, $\mathcal{X}$-Cluster addresses this task by first translating the visual content of the image collection into a lower-dimensional representation, i.e., natural language descriptions. We then exploit the large context window of the LLMs/MLLMs to discover the grouping criteria, followed by their semantic clusters, using the derived textual descriptions.
Concretely, as shown in Figure 3, $\mathcal{X}$-Cluster comprise: (1) a Criteria Proposer that is in charge of discovering the grouping criteria (e.g., Activity, Mood); and (2) a Semantic Grouper that finds semantic clusters by conditioning the predictions on each discovered criterion (e.g., for a given image “Climbing gym” is predicted when conditioned on the criterion Activity, whereas “Adventurous” is predicted when conditioned on Mood). Next, we elaborate on each component of $\mathcal{X}$-Cluster.
4.1 Criteria Proposer
We explore three Proposer designs that convert visual information into text, enabling the LLM to infer grouping criteria. Each design differs in how we convert the visual content into text and in how we use the text to derive the grouping criteria.
Image-based Proposer: We prompt a MLLM (Li et al., 2025) to directly infer grouping criteria $\mathcal{R}$ from images, bypassing the translation of visual content into text. In detail, we present a mini-batch of images, tiled into a composite of $8{\times}8$ image grid, to the MLLM. This allows the MLLM to reason concurrently over a group of images. The image-based approach balances broader coverage of criteria with the preservation of fine-grained details, a trade-off introduced by the tiling operation.
Tag-based Proposer. This variant adopts a tag-driven strategy. We use an open-vocabulary tagger (Radford et al., 2021) to assign up to ten WordNet (Miller, 1995) tags to each image, yielding concise textual descriptors. A LLM (Meta, 2024a) then summarizes all the aggregated tags in $\mathcal{D}$ to propose the grouping criteria $\mathcal{R}$.
Caption-based Proposer (Main). While image tags capture certain visual semantics, they primarily focus on object-related content. To incorporate richer visual cues, such as scene context and interactions, we use a MLLM (Liu et al., 2024b) to caption every image in $\mathcal{D}$. We then split the captions into the largest chunks that fit the LLM’s context window (e.g.128k or 1M tokens) (Meta, 2024a; OpenAI, 2024) and prompt the LLM on each chunk to propose grouping criteria. The deduplicated criteria from all chunks form the pool $\mathcal{R}$.

### v5

@@ Abstract
In this work, we introduce and study the novel task of Open-ended Semantic Multiple Clustering (OpenSMC). Given a large, unstructured image collection, the goal is to automatically discover several, diverse semantic clustering criteria (e.g., Activity or Location) from the images, and subsequently organize them according to the discovered criteria, without requiring any human input. Our framework, $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, treats text as a reasoning proxy: it concurrently scans the entire image collection, proposes candidate criteria in natural language, and groups images into meaningful clusters per criterion. This radically differs from previous works, which either assume predefined clustering criteria or fixed cluster counts. To evaluate $\mathcal{X}$-Cluster, we create two new benchmarks, COCO-4C and Food-4C, each annotated with four distinct grouping criteria and corresponding cluster labels. Experiments show that $\mathcal{X}$-Cluster can effectively reveal meaningful partitions on several datasets. Finally, we use $\mathcal{X}$-Cluster to achieve various real-world applications, including uncovering hidden biases in text-to-image (T2I) generative models and analyzing image virality on social media. Code and datasets will be open-sourced for future research.
Figure 1: Organizing unstructured image collections with $\bm{\mathcal{X}}$-Cluster. Mid: We propose $\mathcal{X}$-Cluster, which takes any unstructured image collection as input, automatically discovers multiple criteria (e.g., Activity or Location) that meaningfully group the data, and outputs images organized into semantic clusters under each criterion, without any prior knowledge. We demonstrate $\mathcal{X}$-Cluster as a versatile tool for real-world analysis: i) Left - uncovers surprising novel biases (e.g., “Hair color” or “Hair style”) when applied to text-to-image (T2I) model outputs, and ii) Right - reveals visual factors (e.g., “Dramatic” color or “Intensive” emotion) that drive social media posts virality.
footnotetext: Corresponding author.
@@ 1 Introduction
When organizing a large collection of unlabelled images, a natural question arises: how should we group them? One could imagine many possible criteria, i.e., based on Activity, Location, or even Color. Yet, it is often unclear which criterion, if any, best describes the dataset, or whether multiple valid grouping principles coexist. As a result, the seemingly simple task of clustering images becomes challenging, as it is influenced by both the visual appearance of the data and their underlying semantics. However, tackling this open-ended unsupervised task of automatically uncovering diverse and interpretable substructures within large image collections is pivotal for many applications, such as social media recommendation [cheng2024retrieval] and dataset auditing [bianchi2023easily].
Existing clustering approaches still heavily rely on an iterative, human-in-the-loop interpretation and refinement process. Typically, we begin by setting a few hyperparameters (e.g., the number of grouping criteria or clusters) for Deep Clustering (DC) methods [caron2018deep, van2020scan] to get a single partition, or for Multiple Clustering (MC) methods [yu2024multiple] to produce several partitions showing different views of the data. We then inspect sample images from each cluster, hoping they correspond to meaningful categories (e.g., “Surfing” or “Skateboarding”) and, ideally, that all clusters follow a coherent criterion (e.g., Activity). When such patterns fail to emerge, we tweak the hyperparameters and try again until the clusters finally make sense to us. This labor-intensive trial-and-error loop exists because i) the resulting clusters are not directly interpretable, being represented only as index assignments. ii) both DC and MC methods converge to solutions shaped by model inductive biases and hyperparameter settings, rather than the data’s intrinsic semantics.
To enhance controllability and interpretability, recent studies have introduced Text-Conditioned Multiple Clustering (TCMC) [yao2024multi, yao2025customized]. TCMC approaches employ Multimodal Large Language Models (MLLMs) [openai2024gpt4o, liu2024llava] to generate semantic clusters based on user-defined criteria and assign images accordingly, producing human-understandable cluster labels. However, these approaches assume that users already know meaningful ways to organize the dataset. As datasets grow in size and complexity, defining such criteria becomes increasingly unrealistic. Moreover, by relying on static preset criteria, this paradigm may overlook previously unknown grouping dimensions that organically emerge from ever-evolving data.
In this paper, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), the first task that aims to automatically generate open-ended and interpretable groupings of large, unstructured image collections without any human priors. Specifically, the goal of OpenSMC is to discover clustering criteria directly from the data and uncover their corresponding semantic clusters to organize images accordingly. This task is particularly challenging because i) it requires concurrent reasoning over all images to identify valid clustering criteria, and ii) it assumes no access to user knowledge about either the clustering criteria or the number of clusters. Tab. 1 summarizes the key differences between OpenSMC and other paradigms.
@@
To address OpenSMC, we make the following contributions. First, we introduce $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a novel training-free two-stage framework powered by MLLMs [liu2024llava, li2024llava, li2023blip] and LLMs [meta2024llama31]. $\mathcal{X}$-Cluster consists of two consecutive modules: the Criteria Proposer and the Semantic Grouper. The Criteria Proposer employs a LLM to holistically reason over the entire image collection through textual representations to discover potential clustering criteria. For each discovered criterion, the Semantic Grouper then organizes images into distinct semantic substructures based on their criterion-related visual content. As shown in Figure 1(mid), our $\mathcal{X}$-Cluster automatically discovers clustering criteria (e.g., Activity, Location) and uncovers their corresponding semantic clusters (e.g., “Surfing”, “Skateboarding” under Activity), all expressed in human-interpretable natural language.
As our second contribution, we introduce two realistic and large-scale benchmarks, COCO-4c and Food-4c, each annotated with ground-truth data for up to four clustering criteria. Using them, we comprehensively evaluate the effectiveness of our method in both criteria discovery and semantic grouping. As our third contribution, we demonstrate the versatility of $\mathcal{X}$-Cluster by applying it across diverse applications. When applied to occupation portrait images generated by text-to-image (T2I) generative models (Figure 1(left)), it uncovers novel occupational biases, such as DALL·E3 [betker2023improving] associating CEOs with “dark” and “short” hair, beyond well-known biases (e.g., Gender). When applied to social media image posts, $\mathcal{X}$-Cluster finds that images featuring “dramatic” colors, “intense” emotions, or ‘urban eclectic” clothing styles tend to attract greater popularity online. These findings show that $\mathcal{X}$-Cluster is a practical tool for understanding large-scale unstructured visual data, enabling the discovery of novel, unexpected patterns.
@@ 2 Related Work
Image Clustering. Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self‑supervision [caron2018deep, caron2020unsupervised, ren2024deep]. Multiple clustering extends this idea, seeking multiple non‑redundant partitions with data augmentations, diversity losses, or subspace methods [qi2009principled, hu2018subspace, yao2023augdmc, metaxas2023divclust, yu2024multiple]. Despite steady progress, both paradigms share key limitations: i) their results are shaped by model inductive biases and training algorithms, limiting generalization beyond object-centric data and often misaligning with user intent or data semantics; and ii) clusters are produced as numeric indices rather than human-readable names. In contrast, $\mathcal{X}$-Cluster derives both meaningful criteria and cluster names directly from unlabeled data.
@@
Topic Discovery. OpenSMC is also related to Topic Discovery [blei2003latent, wang2009multi, eklund2022topic] in NLP, which identifies latent themes [viswanathan2024large, zhang2023clusterllm, zhong2024explaining] or events [tipirneni2024context, nakshatri2023using] from text corpora. Our work similarly aims to uncover common themes from large, unstructured data but operates on images, which is more challenging since i) visual semantics are implicit, unlike text where meaning is explicit, and ii) no current vision model can reliably reason over large image sets.
3 Open-ended Semantic Multiple Clustering
Task Definition. Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering is to build a system, $\mathcal{H}$, that automatically i) discovers a set of $L$ grouping criteria $\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them. Formally, we define an OpenSMC system as:
@@ 3 Open-ended Semantic Multiple Clustering
where each cluster $\mathcal{C}_{k}^{l}$ is characterized by a semantic name $s_{k}^{l}$ and a subset of images $\mathcal{D}_{k}^{l}\subset\mathcal{D}$ that share the same semantics. A criterion $R_{l}$ refers to a theme for grouping images, such that all the clusters under $R_{l}$ should align with the theme. As shown in Figure 1(top), if $R_{l}=\texttt{Activity}$, each cluster under this criterion should collect images $\mathcal{D}_{k}^{l}$ that depict an activity, such as $s_{k}^{l}=\text{``Surfing''}$. If $R_{l}=\texttt{Location}$, the same dataset should be organized into clusters like “Restaurant”, “Sports facility”, and so on.
An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language. In contrast to TCMC setting, where criteria $\mathcal{R}$ and the corresponding cluster counts $K$ are preset by human operators.
Figure 3: $\bm{\mathcal{X}}$-Cluster consists of a Criteria Proposer and a Semantic Grouper. (left) Given a set of images, the Proposer discovers and outputs a pool of grouping criteria in natural language. (right) The Grouper subsequently extracts criterion-specific descriptions from images relevant to each criterion, discovers the underlying semantic clusters, and groups each image at three semantic granularity levels. Results shows an example, as how an unstructured image collection can be grouped into clusters of different semantic granularity corresponding to criterion “Location”. See Supp. F for implementation and prompt details.
Benchmark. Evaluating OpenSMC methods requires benchmarks that can be partitioned under multiple criteria. Currently, only a few benchmarks [yu2024multiple] support the evaluation of OpenSMC methods: Fruit-2c [muresan2018fruit], Card-2c [kaggle2022card], Action-3c [kwon2023image], and Clevr-4c [vaze2024no]. As shown in Figure 2, these benchmarks are limited by their object-centric nature with simple backgrounds (e.g. Fruit-2c), an insufficient number of criteria (e.g. up to three in Action-3c), and a lack of photorealism due to synthetic generation (e.g. Clevr-4c).
Given that the data encountered in real-world applications is more complex, we annotate and propose two new benchmarks for OpenSMC: Food-4c and COCO-4c. Food-4c is sourced from Food-101 [bossard2014food], which includes 101 Food type (original annotations), along with new annotations for 15 Cuisine types, 5 Courses types, and 4 Diet preferences, totaling four clustering criteria. Additionally, we introduced COCO-4c using images from COCO-val [lin2014microsoft], where we annotated four criteria with varying number of clusters: 64 Activity, 19 Location, 20 Mood, and 6 Time of day. Examples of these newly constructed benchmarks are shown in Figure 2. Further details, such as cluster names and the annotation pipeline, are provided in Supp. D.
4 Method
The goal of an OpenSMC system is to first discover meaningful grouping criteria (or themes) from an unstructured image collection by finding commonalities among the images, and then group them into semantic clusters as per the discovered criteria. This is particularly a challenging task because it requires reasoning over the visual content of all images simultaneously. To address OpenSMC, we diverge from representation learning-based MC approaches [yao2024multi, yao2025customized], as no existing model can yet encode large image sets and reason over them reliably. Instead, we convert the visual content of all images into text and use text descriptions as a proxy to discover the grouping criteria and the semantic substructures.
System Overview: As illustrated in Figure 3, our proposed $\mathcal{X}$-Cluster is a two-stage framework that is composed of two modules: Criteria Proposer and Semantic Grouper. The Criteria Proposer processes the entire image set $\mathcal{D}$ to discover diverse common themes among the images and proposes grouping criteria $\mathcal{R}$ in natural language (e.g. Location). Once the criteria are proposed, the Semantic Grouper uncovers the substructure $\mathcal{O}_{l}$ of $\mathcal{D}$ by discovering distinct semantic clusters and assigning images to their respective clusters (e.g. “Tennis Court”), adhering to each criterion $R_{l}\in\mathcal{R}$. As the OpenSMC task operates without user priors to guide semantic granularity, we also design our Semantic Grouper to automatically discovers $\mathcal{O}_{l}$ across multiple granularity levels, from coarse (e.g., “Outdoor”) to mid (e.g., “Recreation”) to fine (e.g., “Tennis Court”), and organizes the images accordingly. In this work, we explore three design choices for both the Proposer and the Grouper. Due to space constraints, only the main variant of $\mathcal{X}$-Cluster is illustrated in Figure 3, while the illustrations of the alternative variants and additional implementation details, including the exact prompts, are provided in Supp. F. Next, we describe each variant in detail.
4.1 Criteria Proposer
As shown in Figure 3(left), the Proposer takes as input a set of input images and generates distinct grouping criteria (or Criteria Pool) in natural language. Its core design principle is the ability to concurrently reason across a large set of images. Next, we explore three systematic approaches.
dlgCaption-based Proposer (main):
To enable reasoning over a large image set for criterion discovery, we first leverage a MLLM [liu2024llava] to generate a comprehensive caption $e_{n}$ for each image, converting its visual content into text representations $e_{n}=\text{MLLM}(\mathbf{x}_{n})$. The resulting caption set $\{e_{n}\}_{n=1}^{N}$ serves as a rich and holistic semantic proxy for the image collection $\mathcal{D}$. Using these compact textual proxies, we then prompt a LLM [meta2024llama31, openai2024gpt4o] to jointly analyze the aggregated visual content and propose multiple valid grouping criteria, denoted as $\widetilde{\mathcal{R}}=\text{LLM}(\{e_{n}\}_{n=1}^{N})$. As an example, the LLM could use different cues such as “Tennis”, “grassy field”, “rock wall” in the captions (see Figure 3) and its reasoning capability to discover the criterion “Location”, since they are usually associated to a physical location, albeit locations may not explicitly appear in any caption.
dlgTag-based Proposer (alternative):
Instead of using captions as textual proxies for reasoning, we further explore an approach that relies on image tags. Specifically, using the WordNet [miller1995wordnet] vocabulary as the candidate tag set, we employ an open-vocabulary tagger (e.g., CLIP [radford2021learning]) to assign ten tags $t_{i,n}$ to each image as $\{t_{i,n}\}_{i=1}^{10}=\text{Tagger}(\mathbf{x}_{n},\text{WordNet})$. These tags act as concise semantic descriptors that summarize the key elements present in each image. We then aggregate all assigned tags and prompt a LLM to analyze them jointly and propose grouping criteria as $\widetilde{\mathcal{R}}=\text{LLM}(\{\{t_{i,n}\}_{i=1}^{10}\}_{n=1}^{N})$.
dlgImage-based Proposer (alternative):
Lastly, we explore an approach that reasons directly over images rather than their textual proxies. Since no existing model can reliably encode large image sets at once, we adopt a simple workaround: divide $\mathcal{D}$ into 64-image batches, stitch each batch into an $8\times 8$ grid as a single composite image, and feed the resulting image grids to a MLLM. The model is prompted to propose grouping criteria from each grid, and we aggregate and deduplicate these subset proposals to obtain the final criteria set $\widetilde{\mathcal{R}}$.